Sally Guthrie is looking for an investment vehicle that will double her money in six years. What interest rate, to the nearest whole percentage, does she have to receive? _______% At that rate, how long will it take the money to triple? Round the answer to 2 decimal places. Do not round your intermediate calculations. ____years If she can't find anything that pays more than 11%, approximately how long will it take to double her investment? Round the answer to 2 decimal places. ____years
Answer #1

1. Let the investment amount be 100 and Interest rate be IR.

Investment Amount + Investment Amount x IR/100 x Time (in years) = 2 x Investment Amount

= 100 + 100 x IR/100 x 6 = 200

= 6IR + 100 = 200\

= 6IR = 100

= IR = 16.67% or 17%

2. Time period for tripling the money invested

Let time required be T

Investment Amount + Investment Amount x IR/100 x Time (in years) = 3 x Investment Amount

= 100 + 100 x 17/100 x T = 300

= 17T = 200

= T = 11.76 or 12 Years

3. If interest rate is not more than 11%, Time required to double the investment

Let time required be T

Investment Amount + Investment Amount x IR/100 x Time (in years) = 2 x Investment Amount

= 100 + 100 x 11/100 x T = 200

= 11T = 100

= T = 9.09 or 9 Years

For any query or clarification, please leave a comment.